Agie is 2 years olderthan Gavin.In 6 years , Gavin's age will be twice Angie's agenow.How old are they 6years from now?
Found 3 solutions by mananth, MathTherapy, ikleyn:
Answer by mananth(16946)

.In 6 years , Gavin's age will be twice Angie's agenow.How old are they 6years from now?
Let Gavin's age be x
Gavin will be x+6 years old 6 years later
Agie is 2 years older than Gavin = x+2
Gavin will be x+2+6 =2x
x=8 that is Gavin's age now
6 years later he will be 14 years
Angie's age now = 8+2 =10 years
After 6 years it will 16 years




Answer by MathTherapy(10552)

Agie is 2 years olderthan Gavin.In 6 years , Gavin's age will be twice Angie's agenow.How old are they 6years from now?
BEWARE..BEWARE...BEWARE Gavin is NOT 8, and NEITHER is Angie 10, as the other person states.
I do wonder when these people will STOP leading others astray with their CONSTANT WRONG answers!!
Subsequently, their ages will 8 and 10, in 6 years' time, and NOT 14 and 16, in 6 years' time, as he/she claims.

Answer by ikleyn(52780)

.
Agie is 2 years older than Gavin. In 6 years, Gavin's age will be twice Angie's age now.
How old are they 6 years from now?
~~~~~~~~~~~~~~~~~~


            To facilitate the solution,  I first edited the post,
            inserting empty space dividers between the words  (as it should be done in any normal text).


Let G be the Gavin's age 6 years from now.

Then his age NOW is (G-6) years.


Agie's age 6 years from now is 2 years more than Gavin's, i.e. (G+2) years.
Hence, Agie's age NOW is ((G+2)-6) = (G-4).


The condition gives us THIS equation

    G = 2*(G-4).

Simplify and solve

    G = 2G - 8

    8 = 2G - G

    G = 8.


ANSWER.  Gavin' age 6 years from now will be 8 years.

         Agie's age 6 years from now will be 8+2 = 10 years.

Solved.